OpenStudy (anonymous):

checkkk my work?(: Identify the oblique asymptote of f(x) = quantity x plus 4 over quantity 3 x squared plus 5 x minus 2 I got y = 3x - 7

OpenStudy (rsadhvika):

check the question again, you didnt flip the expression ?

OpenStudy (rsadhvika):

x+4 ------------ 3x^2+5x-2

OpenStudy (rsadhvika):

its like that ?

OpenStudy (anonymous):

yes

OpenStudy (rsadhvika):

then no oblique asymptote, cuz numerator degree is less. it will have a horizontal asymptote at y=0. thats all

OpenStudy (anonymous):

y = 3x - 7 y = 3x + 17 y = 0 No Oblique Asymptote these are the answer choices..

OpenStudy (rsadhvika):

tick No Oblique Asymptote

OpenStudy (anonymous):

okay. help me with one more?!?!?! pleeeeaseesssss?

OpenStudy (rsadhvika):

sure :)

OpenStudy (anonymous):

Identify the oblique asymptote of f(x) = quantity 4 x squared minus x plus 2 over quantity x plus 1

OpenStudy (anonymous):

and these are my answer choices! y = 4x - 5 y = 4x + 3 y = 0 No Oblique Asymptote

OpenStudy (rsadhvika):

numerator degree > denominator degree so it will have a oblique asymptote when u divide, you would get y = 4x-5

OpenStudy (anonymous):

Thank you so sos so sosososososoo much!

OpenStudy (rsadhvika):

np :)

OpenStudy (lakota1999):

right